# thanks: Authors of equal contributionsthanks: Authors of equal contributions

### v3

@@ III.2 Exactness of fermionic DQME-SQ
$\quad-i\sum_{k}\Big{[}(-)^{M+N-\theta^{+}_{k}}\hat{c}^{-}\rho^{+-}_{{\bf n}_{k}^{+}{\bf m}}-(-)^{\theta^{+}_{k}}\rho^{+-}_{{\bf n}_{k}^{+}{\bf m}}\hat{c}^{-}$
@@ IV Representing DQME-SQ by quantum circuits
It should be mentioned that the unitary decomposition approximation in Eq. (32) will incur a systematic error on the order of $O(\epsilon^{2})$ Sch22023216 . For practical applications on real quantum computers in the noisy intermediate-scale quantum (NISQ) regimes and beyond, these errors and measurement costs from the unitary decomposition approximations can be significant Sch21270503 ; Sch22023216 . A sufficiently large $\epsilon$ is needed to resolve the matrices representing the quantum gates from the device noise, while a small $\epsilon$ is needed to obtain an accurate expansion of Eq. (32). To address this conflict, techniques such as Richardson’s extrapolation can be employed to improve the epsilon error and generate results with higher-order accuracy Sch21270503 ; Sch22023216 . Additionally, amplitude amplification technology can be used to reduce the measurement costs by amplifying the probability of detecting the target state Bra0253 ; Kwo21062438 ; Sur231002 . Therefore, the errors and measurement costs induced by the unitary decomposition approximation in Eq. (32) are thus controllable and expected to be gradually alleviated as quantum computing technologies advance.
Figure 3: (a) Simulated time evolution of $P_{1}-P_{0}$, the population difference between two states, in a spin-boson model, by using the quantum algorithm for the DQME-SQ theory, compared to low and high temperature results via the classical propagation of HEOM; (b) Simulated time evolution of the ground-state population $P_{0}$ for a single-impurity Anderson model, with inset depicting the imaginary part of the correlation function. Detailed parameters are provided in the Appendix C.
@@ V Numerical demonstration through digital quantum simulation
Simulations of more complex OQSs are demonstrated in the Appendix C. These additional examples further exhibit the exactness and universality of the DQME-SQ theory. Due to limited computational resources at our disposal, the digital quantum simulations presented above are constrained to minimal models. When implementing the DQME-SQ approach on realistic NISQ devices, the errors and measurement costs arising from the unitary decomposition approximations will need to be carefully considered and addressed. Techniques such as those proposed in Refs. Sch21270503 ; Sch22023216 can be employed to mitigate these issues. However, as quantum computing technologies continue to advance significantly in the future, the utilized quantum algorithm, including the LCU-Trotter method, will become universally feasible for simulating more complex OQSs.
@@ Appendix A Relationship between bosonic DQME-SQ and pseudomode theories
$\dot{\rho}_{\bf{u,v}}(t)=$ $-i[H_{\mbox{\tiny S}},\rho_{\bf{u,v}}]-\sum_{k}\gamma_{k}\Big{(}u_{k}\rho_{{\bf u}_{k}^{-},{\bf v}_{k}^{+}}+v_{k}\rho_{{\bf u}_{k}^{+},{\bf v}_{k}^{-}}\Big{)}$
@@ Appendix B Trace-preserving property of reduced density operator in DQME-SQ
Appendix C Details on the numerical implementation of DQME-SQ
@@ 2. Examples on more complex OQSs
The related parameters are given by $W=50\Delta$, $U=U_{\rm C}=12\Delta$, $t=10\Delta$, and temperature $T=5\Delta$. The number of dissipaton levels is chosen to be $4$ ($K=1$ for each spin orbital), using the classical HEOM and DEOM-SQ algorithms. All the dissipaton configurations are included without any truncation. As illustrated in Fig. 5, the results from these two methods match perfectly, where $P_{1}\equiv\langle 1_{1\uparrow}1_{1\downarrow}0_{2\uparrow}0_{2\downarrow}|\rho_{\mbox{\tiny S}}|1_{1\uparrow}1_{1\downarrow}0_{2\uparrow}0_{2\downarrow}\rangle$ and $P_{\text{double}}\equiv\langle 1_{1\uparrow}1_{1\downarrow}1_{2\uparrow}1_{2\downarrow}|\rho_{\mbox{\tiny S}}|1_{1\uparrow}1_{1\downarrow}1_{2\uparrow}1_{2\downarrow}\rangle$.

### v4

@@ III.2 Exactness of fermionic DQME-SQ
$\quad-i\sum_{k}\Big{[}(-)^{M+N-\theta^{+}_{k}}\hat{c}^{-}\rho_{{\bf n}_{k}^{+}{\bf m}}-(-)^{\theta^{+}_{k}}\rho_{{\bf n}_{k}^{+}{\bf m}}\hat{c}^{-}$
@@ IV Representing DQME-SQ by quantum circuits
It should be mentioned that the unitary decomposition approximation in Eq. (32) will incur a systematic error on the order of $O(\epsilon^{2})$ Sch22023216 . For practical applications on real quantum computers in the noisy intermediate-scale quantum (NISQ) regimes and beyond, these errors and measurement costs from the unitary decomposition approximations can be significant Sch21270503 ; Sch22023216 . A sufficiently large $\epsilon$ is needed to resolve the matrices representing the quantum gates from the device noise, while a small $\epsilon$ is needed to obtain an accurate expansion of Eq. (32). To address this conflict, techniques such as Richardson’s extrapolation can be employed to improve the epsilon error and generate results with higher-order accuracy Sch21270503 ; Sch22023216 . Additionally, amplitude amplification technology can be used to reduce the measurement costs by amplifying the probability of detecting the target state Bra0253 ; Kwo21062438 ; Ber; Sur231002 . After considering both error control and amplitude amplification, the circuit depth of the LCU-Trotter algorithm grows polynomially as the degrees of freedom of dissipaton increases. The detailed analysis of complexity and error is presented in Appendix C. Therefore, the errors and measurement costs induced by the unitary decomposition approximation in Eq. (32) are thus controllable and expected to be gradually alleviated as quantum computing technologies advance.
Figure 3: (a) Simulated time evolution of $P_{1}-P_{0}$, the population difference between two states, in a spin-boson model, by using the quantum algorithm for the DQME-SQ theory, compared to low and high temperature results via the classical propagation of HEOM; (b) Simulated time evolution of the ground-state population $P_{0}$ for a single-impurity Anderson model, with inset depicting the imaginary part of the correlation function. Detailed parameters are provided in the Appendix D.
@@ V Numerical demonstration through digital quantum simulation
Simulations of more complex OQSs are demonstrated in the Appendix D. These additional examples further exhibit the exactness and universality of the DQME-SQ theory. Due to limited computational resources at our disposal, the digital quantum simulations presented above are constrained to minimal models. When implementing the DQME-SQ approach on realistic NISQ devices, the errors and measurement costs arising from the unitary decomposition approximations will need to be carefully considered and addressed. Techniques such as those proposed in Refs. Sch21270503 ; Sch22023216 can be employed to mitigate these issues. However, as quantum computing technologies continue to advance significantly in the future, the utilized quantum algorithm, including the LCU-Trotter method, will become universally feasible for simulating more complex OQSs.
@@ Appendix A Relationship between bosonic DQME-SQ and pseudomode theories
$\dot{\rho}_{\bf{u,v}}(t)=$ $-i[H_{\mbox{\tiny S}},\rho_{\bf{u,v}}]-\sum_{k}\gamma_{k}\Big{(}u_{k}\rho_{{\bf u},{\bf v}}+v_{k}\rho_{{\bf u},{\bf v}}\Big{)}$
@@ Appendix B Trace-preserving property of reduced density operator in DQME-SQ
Appendix C Detailed analysis of error and complexity
In this appendix, we will present more detailed analysis of error and complexity for DQME-SQ-based LCU-Trotter algorithm.
The LCU-Trotter method exhibits an error of order $O(\epsilon^{2})+O(\lambda K{\Delta}t^{2})$ for evolution over a time step ${\Delta}t$. Here, $K$ represents the number of dissipaton modes, and $\lambda$ characterizes the system-environment coupling strength, with the dimension of $[T]^{2}$. The goal is to control the error within the precision, Err, during the evolution from time 00 to $t$. Given that dissipative dynamics include decay terms that suppress divergence, it is assumed that the errors in each time step during the evolution are linearly additive. Thus, the total error is approximately $tO(\epsilon^{2}/{\Delta}t)+tO(\lambda K{\Delta}t)\sim\text{Err}$. This leads to the requirement that the step length be constrained as ${\Delta}t\sim\text{Err}/(t\lambda K)$, and the expansion coefficient as $\epsilon\sim\text{Err}/(t\lambda^{1/2}K^{1/2})$. Under this error control, as the number of dissipaton modes $K$ increases, the step size ${\Delta}t$ decreases. The circuit depth grows polynomially ($\sim O(\lambda^{1/2}t^{2}\text{Err}^{-1}K^{2})$) as the number of dissipaton modes increasing.
After a single step of evolution, the amplitude of the target state $|\tilde{\rho}\rangle\otimes|0\rangle$ is $2\epsilon$, so the probability of measuring the target state is only $4\epsilon^{2}$ after performing a Z-measurement on the auxiliary qubit. Since the target state $|\tilde{\rho}\rangle\otimes|0\rangle$ is orthogonal to the unwanted state $|\text{unwanted}\rangle\otimes|1\rangle$, this issue can be addressed using existing amplitude amplification techniques Bra0253 ; Ber14283 ; Kwo21062438 ; Sur231002 to increase the measurement probability to the order of 1. Amplitude amplification, which shares its origins with the Grover search algorithm, works by applying multiple rotations to the superposition of the target state and its orthogonal state, thus enhancing the target state’s amplitude and improving the measurement probability. The number of rotations depends on $\epsilon$, scaling as $O(\epsilon^{-1})$. The circuit depth for each rotation operation is proportional to the LCU-Trotter circuit depth, which is $O(K)$ Ber14283 . In conclusion, after considering both error control and amplitude amplification, as the number of dissipaton modes $K$ increases, the circuit depth of the LCU-Trotter algorithm grows polynomially ($\sim O(\lambda^{3/2}t^{3}\text{Err}^{-2}K^{5/2})$). There is still room for further optimization in this regard. In the future, we also aim to combine the DQME-SQ framework with more advanced quantum algorithms to simulate complex open quantum systems more efficiently and accurately.
Appendix D Details on the numerical implementation of DQME-SQ
@@ 2. Examples on more complex OQSs
The related parameters are given by $W=50\Delta$, $U=U_{\rm C}=12\Delta$, $t=10\Delta$, and temperature $T=5\Delta$. The number of dissipaton levels is chosen to be $4$ ($K=1$ for each spin orbital), using the classical HEOM and DQME-SQ algorithms. All the dissipaton configurations are included without any truncation. As illustrated in Fig. 5, the results from these two methods match perfectly, where $P_{1}\equiv\langle 1_{1\uparrow}1_{1\downarrow}0_{2\uparrow}0_{2\downarrow}|\rho_{\mbox{\tiny S}}|1_{1\uparrow}1_{1\downarrow}0_{2\uparrow}0_{2\downarrow}\rangle$ and $P_{\text{double}}\equiv\langle 1_{1\uparrow}1_{1\downarrow}1_{2\uparrow}1_{2\downarrow}|\rho_{\mbox{\tiny S}}|1_{1\uparrow}1_{1\downarrow}1_{2\uparrow}1_{2\downarrow}\rangle$.